a scientist has 2 solutions, which he has labeled solituion A and solutions B. each contain salt. he knows that solution A has 60% salt and solutions B has 80% salt. he wants to obtain 40 ounces of a mixture that is 75% salt. how many ounces of each solution should he use?
______ounces of solution A
______ ounces of solutions B

Answer by lwsshak3(11628)

a scientist has 2 solutions, which he has labeled solituion A and solutions B. each contain salt. he knows that solution A has 60% salt and solutions B has 80% salt. he wants to obtain 40 ounces of a mixture that is 75% salt. how many ounces of each solution should he use?
***
let x=amt of solution A to use
40-x=amt of solution B to use
..
60%x+80%(40-x)=75%*40
.60x+32-.80x=30
0.20x=2
x=10
40-x=30
amt of solution A to use= 10 ounces
amt of solution B to use= 30 ounces